OpenStudy (anonymous):

A flower bed is in the shape of a triangle, with a base of 13 feet and a height of 5 feet. What is the area of the flower bed? a.18 square feet b.32.5 square feet c.65 square feet d.130.5 square feet

OpenStudy (anonymous):

@anna_truong123 @quickstudent @wil476003 @Ibbutibbu.

OpenStudy (anonymous):

The area for a triangle is A=1/2bh so it's (13)(5)/2 = b.32.5

OpenStudy (mathstudent55):

base times height divided by 2

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

\(A = \dfrac{bh}{2} = \dfrac{13~ft \times 5~ft}{2} = \dfrac{65~ft^2}{2} = 32.5~ft^2\)